OpenStudy (goformit100):

In a cricket match, five batsmen A, B, C, D and E scored an average of 36 runs. D Scored 5 more than E; E scored 8 fewer than A; B scored as many as D and E combined; and B and C scored 107 between them. How many runs did E score ? A. 62 B. 45 C. 28 D. 20

OpenStudy (goformit100):

@ash2326 @Preetha

OpenStudy (anonymous):

I think it's D. 20 This is how I figured it out, Total runs scored = (36 x 5) = 180. Let the runs scored by E be x. Then, runs scored by D = x + 5; runs scored by A = x + 8; runs scored by B = x + x + 5 = 2x + 5; runs scored by C = (107 - B) = 107 - (2x + 5) = 102 - 2x. So, total runs = (x + 8) + (2x + 5) + (102 - 2x) + (x + 5) + x = 3x + 120. Therefore 3x + 120 =180 => 3X = 60 => x = 20